a rectangular garden measures 20 ft by 32 ft. Surrounding the garden is a path 2ft wide. What is the area of the path?
Answer by stanbon(75887)

a rectangular garden measures 20 ft by 32 ft. Surrounding the garden is a path 2ft wide. What is the area of the path?
Draw a rectangle inside a larger rectangle.
----
Area of the inner rectangle = 20*32 = 640 sq ft
------
Area of the outer rectangle = (20+4)(32+4) = 24*36 = 144*6 = 864 sq. ft
----
path area = 864-640 = 224 sq ft.
----------------------
Cheers,
Stan H.
-------------